OpenStudy (anonymous):

choose the replacement for x that makes the equation true -32x+2=-62

OpenStudy (anonymous):

-32x = 62 - 2 x = -62 - 2 + 32 x ?

OpenStudy (anonymous):

-32? .... thats not one of my options.